# shapes, arrows.meta, positioning Chorus: Foundation Models for Unified Data Discovery and Exploration

### v1

@@ 4. Experiments
DoDuo provides two embedding variants: one trained on the WikiTables dataset and another on VizNet. When using DoDuo as a baseline we test against both, labelling them DoDuo-Wiki and DoDuo-Viz respectively.
| Dataset Title | # Tables | Avg. # Columns | Avg. Rows |
| T2d-class v2 | 237 | 7.41 | 118 |
| VizNet | 3.03 | 5 200 | |
| GitNotebooks | 300 | 10.3 | 78 |
| Overall | 867 | 6.67 | 2 038 |
Datasets
Table 1 outlines the three experiment benchmarks we use. For the table-class annotation task, the T2Dv2 dataset (Ritze and Bizer, 2017), a “gold standard” corpus of approximately one thousand tables, manually annotated by experts. Of those, are annotated with one of DBPedia.org classes. These tables were in turn selected from the Common Crawl corpus of web tables (Foundation, 2011). For column-type annotation, we sample a subset of the VizNet dataset (Hu et al., 2019), extracted by the Sherlock team (Hulsebos et al., 2019) comprising 1 000 columns from approximately 330 tables. For the join-column prediction task, we use a dataset we call GitNotebooks, extracted by the Auto-suggest team (Yan and He, 2020). We select 300 tables from that dataset for which we have join data. For the first two tasks, which require defining a type system for classes and properties, we use the DBPedia ontology (Mendes et al., 2012) for our experiments. This is a community-sourced ontology and is the standard in previous studies.
Setup
We use the GPT-3.5 model (Ouyang et al., 2022) as it is the most widely-available large model with api access at the time of writing. All other code was run on a commodity laptop with 8 physical arm cores and 16GB of main memory. Running all experiments, including development, came to $2 in api costs.
We evaluate using the metrics precision, recall and score. Precision is the proportion of true positive results out of the total predicted positive results, while recall is the proportion of true positive results out of the total actual positive results in the dataset. The score is the harmonic mean of precision and recall. Since we deal with a multiclass setting, we calculate these metrics for each class separately then aggregate by taking the mean, weighted by the class size. Class-weighted precision, recall and are the standard metrics in prior work (Zhang et al., 2020; Suhara et al., 2022; Hulsebos et al., 2019; Cafarella et al., 2008).
4.1. Table-class detection
In this task, we tag each table with the DBPedia ontology entry that represents the row-type of the data. Of the datasets that comprise the T2Dv2 dataset, tables have table-class correspondences available while 763 do not. We denote this subset T2D-class v2. We note that only classes are utilized in this “gold standard” mapping, while DBPedia ontology has classes.
We compare against the baselines DoDuo and TaBert. No approach in the prior work provides out-of-the-box capabilities on this task, so add a classification layer on top of the embedding layer. After computing the column embeddings, predictions are extracted by adding a pooling layer, fed to a multi-layer perceptron, and then finally taking the soft-max. This is a straightforward method of adapting the embeddings to our multi-class setting, used in prior benchmarks for table-class detection (Koleva et al., 2021). We fix the embeddings to their pretrained values and learn the weights of the classification layer using five-fold cross-validation.
Supervised variant
To allow for comparisons with prior work, we initially restrict our system to picking out of the classes. This is because all other approaches require training on labelled instances—the baselines cannot predict outside those classes. We test classes rather than because the classes that occur only once cannot be meaningfully predicted by the baselines that require training labels.
Table 2 shows the results. Chorus improves on the three baselines—DoDuo-Viz, DoDuo-Wiki and TaBERT—on all metrics. score is improved by 0.169 points, precision by percentage points and recall by percentage points. Of the baselines, DoDuo-Wiki provides the best and precision, while TaBERT provides the comparable recall. The best performing models, TaBERT and DoDuo-Wiki are trained on CommonCrawl, a superset of the T2Dv2 benchmark. DoDuo-Viz which is trained on the VizNet, a dataset disjoint from T2Dv2, has the weakest performance. The numbers for TaBERT are in line with prior replications (Koleva et al., 2021), while to the best of our knowledge this is the first benchmarking of DoDuo on this task.
Unsupervised variant
Next, we relax the classification domain, allowing the foundation model to choose any class in the ontology. We then compare the quality of the classes with that of the human-expert labels.
For of tables, our system produces correct results. Of that portion, percentage points are comprised of exact matches, while percentage points are better-than-correct results. This means the predicted labels are clearly and unambiguously better than those selected by the expert. Because this is a strong claim, we list all such datasets in Table 3.
For the final the answer is incorrect: this can mean the answer is completely wrong or simply worse than the label provided by the expert. This means that on the relations where chorus and the expert-label disagree, our system is more likely to be correct.
| Table-class correctness | |||
| -score | Precision | Recall | |
| DoDuo-Viz | 0.654 | 66.8% | 68.3% |
| DoDuo-Wiki | 0.757 | 78.6% | 76.9% |
| TaBERT | 0.746 | 76.3% | 76.8% |
| Chorus | 0.926 | 96.1% | 92.4% |
| ID | Predicted label | Expert label | Evidence and explanation |
Military-Person |
Monarch |
Table includes the values “emperor,” “general” and “rebel” |
Roman-Emperor |
Monarch |
Page titled Crococalana History, a Roman settlement |
Temple |
Monarch |
Text on page: “Das sind vor allem ägyptische Tempel” translates to “These are mainly Egyptian temples” |
Dam |
Lake |
Table is from swiss-dams.ch |
Roman-Emperor |
Monarch |
Page description is Roman Rulers |
Sky-scraper |
Building |
Page titled The Tallest Buildings |
Person |
Scientist |
Famous figures including “Walt Disney” |
Nobel-Prize |
Scientist |
Chemistry Nobel page has one row for each year, including “Not awarded this year” |
Currency |
Country |
Table titled Currency Exchange Rates |
Political-Party |
Election |
Rows of British Elections 2014 are political parties including “Labor” and “Tories” |
Dam |
Lake |
Rows exclusively river dams e.g. “Tinaroo Falls Dam” |
Song |
Work |
From url .../SongList.asp |
Fictional-Character |
Tele-vision-Show |
Rows for “Ryan Atwood” and “Caleb Nicol”, characters from the TV show The O.C. |
GolfTour-nament |
Golf-Player |
Column headers are “Year”, “Winner” and “Second Place” |
Roman-Emperor |
Monarch |
Page mentions ancient Rome, headers are “Emperor”, “reign” |
Roman-Emperor |
Monarch |
url: .../Rome/Emperors.html |
Mathe-matician |
Scientist |
Table from the Mathematics Genealogy Project |
Nobel-Prize |
Scientist |
Nobel prizes list includes columns “Year”, “Winner1”, “Winner2” and “Winner3” |
Political-Party |
Election |
Rows include “Labour Party” and “Conservative Party” |
BodyOf-Water |
Lake |
Values include “Moose River”, “Loon Stream” and “Alder Pond” |
Religious-Organization |
Country |
Text on page: “The table above is generated […] for each diocese” |
4.2. Column-type annotation
Next, we compare the ability of our system to assign classes to table columns.
VizNet is collection of tables, extracted by the Sherlock (Hulsebos et al., 2019) team from the VizNet repository (Hu et al., 2019) of data visualizations and open datasets. VizNet comprises 31 million datasets in total. We selected challenging and mutually exclusive DBPedia.org classes to test: artist, address, city, director, album, currency, product, isbn, manufacturer, birthday. We then used stratified sampling to select 1 000 columns of these types to predict.
Baselines
We compare against TaBERT (Yin et al., 2020), DoDuo (Suhara et al., 2022) and Sherlock (Hulsebos et al., 2019) on this task. Since Sherlock is designed for column annotation, we use the out-of-the-box model provided by the original team. We restrict Sherlock to the ten target classes. For TaBERT we train an additional classification layer on top of the pre-trained embeddings that these frameworks provide. DoDuo provides a classification layer, however unlike Sherlock the current api does not expose raw probabilities, so we cannot restrict it to the target classes. Instead, we chose to add a multilayer perceptron classification layer, as for TaBERT, to give DoDuo fair odds. We fix the embeddings to their pretrained values and learn the weights of the classification layer using five-fold cross-validation.
Results
Table 4 contains the results for the VizNet dataset. Our fm-based approach improves performance on the measured metrics of -score, precision and recall. The best performing method is Sherlock, narrowly beating DoDuo-VizNet, with a 0.930 score. If we consider methods which are not specifically pretrained on VizNet (recall, which is also the test set) chorus is the best performing on all three metrics. It has comparable and precision to Sherlock, but percentage points lower recall.
Note in particular DoDuo-Wiki, which does not have access to VizNet at pretraining time, has a large regression in performance compared to DoDuo-Viznet, losing 0.085 points. We sanity-checked the low scores of TaBERT by replicating its previously reported column-type annotation scores from prior work and were successful.
| -score | Precision | Recall | |
| DoDuo-VizNet | 0.900 | 90.3% | 89.9% |
| Sherlock | 0.930 | 92.2% | 93.1% |
| TaBERT | 0.380 | 38.9% | 38.3% |
| DoDuo-Wiki | 0.815 | 82.6% | 81.4% |
| Chorus | 0.865 | 90.1% | 86.7% |
4.3. Join-column Prediction
Finally, we evaluate our approach’s ability to suggest which columns are the correct choice for a join, the join-column prediction task. We use the GitNotebooks dataset from (Yan and He, 2020), a collection of 4 million Python notebooks (and their associated relational tables) including 24 thousand joins collected from Github. As one of the baslines, Trifacta Wrangler, requires manual execution and recording of each prediction, we restrict this benchmark to 300 randomly sampled tables.
Baselines
For this task, we compare with three baselines. Jaccard similarity, , is the first. Two columns are selected such that where . This is a commonly used approach in the literature (Chen et al., 2014; Dasu et al., 2002; Nargesian et al., 2023; Yan and He, 2020). Another baseline is Levenshtein distance (Levenshtein, 1966), which selects the pair of column names with the smallest edit distance between them. The final baseline is Trifacta Wrangler (Trifacta, 2023), a commercial product spun off from the Wrangler research line (Kandel et al., 2011). When joining two tables in this product, it suggests the keys on which to join them. As no api was available, we obtain all Trifacta predictions by running the joins manually.
Results
Table 5 shows the quality of estimates for our approach and the baselines. We measure the quality of the predictions by the same criteria as the previous tasks. By these metrics, our approach improves the quality of predictions and beats the next-best approach by a clear margin: score is improved by , precision by percentage points and recall by percentage points.
| -score | Precision | Recall | |
| Jaccard | 0.575 | 60.7% | 54.7% |
| Levenshtein | 0.718 | 72.3% | 71.3% |
| Trifacta Wrangler | 0.823 | 82.6% | 82.0% |
| Chorus | 0.895 | 91.0% | 88.0% |
4.4. Dataset contamination
Here we perform an experiment to validate whether any of the testing data occurred in the training corpus of the large-language model, an issue called dataset contamination or data leakage. Because these models are trained on terabytes of internet data (Gao et al., 2021) and we use public benchmarks, they may have seen the test data in training.
We test on seven guaranteed-unseen tables and their columns, all uploaded between April–June 2023 to the federal data repository Data.gov. Since the foundation model training was completed on or before March 2023, these datasets cannot have contaminated its corpus. Repeating the supervised column-type annotation task as in Section 4.2, we measure a 0.857 score, 90.0% precision and 81.8% recall. This is within points, 0.1% precision and 5% recall of the benchmark results.
5. Discussion
Our experiments show that chorus is a very promising approach for data-discovery tasks. The system is able to provide competitive performance on all three tested tasks. For 1⃝ table-class detection and 3⃝ join-column detection, it surpasses all the baselines by a clear margin. On 2⃝ it performs best of the methods not pretrained specifically on the test dataset and comparably with those which are.
Importantly, the performance is robust: it consistently performs well, unlike other baseline methods. For example, TaBERT is the closest to the best performing baseline on T2Dv2 but the worst baseline on VizNet. Similarly DoDuo-Viz performs well on VizNet but loses 0.11 points when applied to T2Dv2 compared to DoDuo-Wiki. Conversely, DoDuo-Wiki loses 0.09 points compared to DoDuo-Viz when tested on the VizNet-based task.
Training data collection
A major advantage of a foundation-model approach is that extensive domain-specific training data need not be collected. In particular, representation learning approaches require large amounts of data for learning the embeddings as well as data for learning the task (in some cases this can be the same dataset). For example, TaBERT requires 26 million tables for training its embeddings. This cost may be acceptable if the embeddings generalize—in paragraph out-of-domain performance we argue they are not. Even with the embeddings fixed, lots of task-specific labels are needed. In (Deng et al., 2020), the use of 250 labels for one task is considered a “small dataset” by the authors and leads to subpar performance. In contrast, our prompts in Figure 5 use one example each for the table-class detection and column-type annotation and zero examples for the join-column prediction task.
Out-of-domain performance
We note a troubling pattern of a lack of cross-domain generalization in representation-learning approaches. In particular, the tested baselines degrade when used to embed tables not from the dataset the embeddings were trained on. Despite being described as “pretrained”, this is in contrast to word embeddings such as word2vec, which perform consistently across a number of data distributions. This finding is in-line with prior work: regressions of up to 0.40 points when switching to new datasets have been observed by other researchers (Deng et al., 2020, Table 9).
This is made worse by the computational cost of training the embeddings: tabert was trained using a cluster of 128 of Nvidia’s Tesla V100 gpus (Yin et al., 2020, A.2), at the suggested retail price of $14 000 each (Corp, 2018). The cost of training fms is even greater, of course, but need only be borne once by the fm developer—not the end user.
Flexibility
Another advantage of chorus we observe in the experiments is task adaptability. In the 1⃝ table-class detection task, we are able to switch the prediction domain easily. Restricting to the 33 classes used by the benchmark can be done by providing the permitted classes to the foundation model; allowing the model to generalize to other DBPedia classes (the unsupervised heading of Section 4.1) is as simple as omitting those instructions. Contextual information, such as table title or url, could be as easily added. In the representation learning setting, such modifications would require retraining the embeddings.
Limitations and risks
We control the risk of dataset contamination by testing for it in Section 4.4. We find that the performance of chorus on guaranteed-unseen datasets is comparable to those in public benchmarks, so good performance on the those benchmarks cannot be explained away as simple data contamination. Seperately, formal linguistic fluency means that errors may fool human reviewers. It is well known that linguistic fluency strongly influences the perception of correctness and quality (Langford, 1999). As such, foundation models may be able to obtain favorable evaluation from testers by producing answers which are plausible but incorrect. This has been called subtle misinformation in prior work (Rae et al., 2021). This is particularly relevant to the evaluation of Table 3. Finally, we note that there can be a large variance in result quality caused by minor changes in the prompt (Zhao et al., 2021).
5.1. Future directions
Additional tasks
The promising performance on our three tasks may extend to many more. Other tasks related to data discovery and exploration that could be explored include schema auto-completion (Cafarella et al., 2008), where missing parts of a partial schema are suggested to the user; join-graph traversal, where successive tables to join on are suggested (Dong et al., 2022); and outlier detection, where potential erroneous data are detected.
However, take care when selecting tasks. Foundation models are not a panacea. For example, they virtually never recall facts that occur less than 10-100 times in the corpus (Kandpal et al., 2022). This means that point lookups of data are unlikely to be supported by unassissted foundation models, at least for the foreseeable future. Further, we emphasize that many foundation models are trained less than one epoch, meaning they see each token in the corpus one time or not at all.
Private or domain-specific datasets
As with all tested baselines, the foundation models are trained on public data. The distribution of data in the public sphere significantly defers from that in specialized domains or private data. Whether observed capabilities hold on e.g. enterprise data lakes is worth investigating. Further application to domain-specific ontology such as dron, a pharmaceutical ontology of drugs only, could be investigated.
Larger and smaller models
Larger models are of interest because if scaling laws continue to hold, their performance should improve (Wei et al., 2022b). The larger GPT-4 model has been announced (Bubeck et al., 2023). However, at the time of experiments the authors did not have API access to this model, so we leave exploring its capabilities to future work. On the other hand, the large cost of the current generation of foundation models has sparked interest in model distillation: reducing the size of the model without incurring a reduction in capabilities (West et al., 2022; Dettmers et al., 2023). However, these downsized models have significantly degraded emergent abilities (Gudibande et al., 2023) and so would likely perform worse on precisely the tasks of interest to us.
End-to-end evaluation
With foundation-model based data discovery systems becoming closer to reality, a holistic evaluation would be valuable to evaluate the utility of this approach to data analysts. This is important because many benchmarks are limited to sparsely-labelled portions of publicly available datasets.
6. Related Work
@@
| Trifacta Wrangler (Trifacta, 2023) | ✗ | ✗ | ✓ |
@@
⚫ requires modification and training data collection

### v3

@@ 4. Experiments
| Trifacta Wrangler (Trifacta, 2023) | ✗ | ⚫ | ✓ |
@@
⚫ required modification or training data collection (see text)
DoDuo provides two embedding variants: one trained on the WikiTables dataset and another on VizNet. We label them DoDuo-Wiki and DoDuo-Viz.
| Dataset Title | # Tables | Avg. # Columns | Avg. Rows |
| T2d-class v2 | 237 | 7.41 | 118 |
| VizNet | 3.03 | 5 200 | |
| GitNotebooks | 24 579 | 30.9 | 60 242 |
| Overall | 35 416 | 23.0 | 43 491 |
Datasets
Table 2 outlines the three experiment benchmarks we use. For the table-class detection task, we test on the T2D-class v2 dataset (Ritze and Bizer, 2017), a “gold standard” corpus of tables, manually annotated by experts with one of DBPedia.org classes. These tables were in turn selected from the Common Crawl corpus of web tables (Foundation, 2011). For column-type annotation, we sample a subset of the VizNet dataset (Hu et al., 2019), extracted by the Sherlock team (Hulsebos et al., 2019), comprised of columns with one of fifteen types from approximately tables. This is in line with prior work that uses VizNet (Hulsebos et al., 2021). For the join-column prediction task, we use a dataset we call GitNotebooks, extracted by the Auto-suggest team (Yan and He, 2020). We select 300 tables from that dataset for which we have join data to run manually. Here we use a sample as one of the baselines, Trifacta Wrangler, does not have an API but instead predictions must be produced manually. Separately, we run all 24 thousand tables on the baselines with an api. For the first two tasks, which require defining a type system for classes and properties, we use the DBPedia ontology (Mendes et al., 2012) for our experiments. This is a community-sourced ontology and is the standard in previous studies.
Setup
We use the GPT-3.5 model (Ouyang et al., 2022) as it is the most widely-available large model with api access at the time of writing. All other code was run on a commodity laptop with 8 physical arm cores and 16GB of main memory. Running all experiments came to a total of $20 in api costs.
We evaluate using the metrics precision, recall and score. Precision is the proportion of true positive results out of the total predicted positive results, while recall is the proportion of true positive results out of the total actual positive results in the dataset. The score is the harmonic mean of precision and recall. Since we deal with a multiclass setting, we calculate these metrics for each class separately then aggregate by taking the mean, weighted by the class size. Weighted precision, recall and are the standard metrics in prior work (Zhang et al., 2020; Suhara et al., 2022; Hulsebos et al., 2019; Cafarella et al., 2008). We also report average throughput and cost for each task.
4.1. Table-class detection
For the first task, 1⃝ table-class detection, we tag each table with the DBPedia ontology entry that represents the row-type of the data. Of the datasets that comprise the T2Dv2 dataset, tables have table-class correspondences available while 763 do not—we exclude the unlabelled ones from the supervised evaluation. We call this subset of annotated tables T2D-class v2 and use it for evaluation on this task. We note that only classes are utilized in this “gold standard” mapping, while DBPedia ontology has classes.
We compare against the baselines DoDuo and TaBert. No approach in the prior work provides out-of-the-box capabilities on this task, so we add a classification layer on top of the pretrained embedding layer. After computing the column embeddings using DoDuo or TaBert, predictions are extracted by adding a pooling layer, fed to a multi-layer perceptron, and then finally taking the soft-max. This is a straightforward method of adapting the embeddings to our multi-class setting, used in prior benchmarks for table-class detection (Koleva et al., 2021). We fix the embeddings to their pretrained values and learn the weights of the classification layer using five-fold cross-validation.
Supervised variant
To allow for comparisons with prior work, we initially restrict our system to picking out of the classes. This is because all other approaches require training on labelled instances—the baselines cannot predict outside those classes. We test classes rather than because the classes that occur only once cannot be tested on baselines that require supervised training (DoDuo and TaBERT), since a result requires a disjoint training and test set.
Table 3 shows the results. Chorus improves on the three baselines on all metrics. score is improved by 0.169 points, precision by percentage points and recall by percentage points. Of the baselines, DoDuo-Wiki provides the best and precision, while TaBERT provides the comparable recall. The best performing models, TaBERT and DoDuo-Wiki are trained on CommonCrawl, a superset of the T2Dv2 benchmark. DoDuo-Viz which is trained on the VizNet, a dataset disjoint from T2Dv2, has the weakest performance. The numbers for TaBERT are in line with prior replications (Koleva et al., 2021), while to the best of our knowledge this is the first benchmarking of DoDuo on this task.
Unsupervised variant
Next, we relax the classification domain, allowing the foundation model to choose any of the classes of the DBPedia ontology. We then compare the quality of the classes with that of the human-expert labels. DoDuo and TaBERT are not evaluated in this task setting as they cannot predict outside the classes they have observed in training.
For of tables, our system produces correct results. Of that portion, percentage points are comprised of exact matches, while percentage points are better-than-correct results. This means we judge the predicted labels are clearly and unambiguously better than those selected by the benchmark authors. This is a strong claim so we list all such datasets in the technical report (Kayali et al., 2023), with evidence. For the final the answer is incorrect: this can mean the answer is wrong or simply worse than the label provided by the expert. This means that on the relations where chorus and the expert-label disagree, our system is more likely to be correct.
Chorus has a throughput of nearly 31 tables per second on this benchmark and cost an average 2.5¢ per 100 table-class predictions.
| -score | Precision | Recall | |
| DoDuo-Viz | 0.654 | 66.8% | 68.3% |
| DoDuo-Wiki | 0.757 | 78.6% | 76.9% |
| TaBERT | 0.746 | 76.3% | 76.8% |
| Chorus | 0.926 | 96.1% | 92.4% |
4.2. Column-type annotation
Next, we compare the ability of our system to assign classes to table columns. VizNet is a collection of tables, extracted by the Sherlock (Hulsebos et al., 2019) team from the VizNet repository (Hu et al., 2019) of data visualizations and open datasets. VizNet comprises 31 million columns in total, of which a test set of 142 000 can be used for evaluation—the rest have trained on by DoDuo and Sherlock. Of those, we select a subset of classes which are supported by both DoDuo-Wiki and DoDuo-VizNet (these baselines support a disjoint set of classes), arriving at test columns used as a benchmark in this section.
Baselines
We compare against TaBERT (Yin et al., 2020), DoDuo (Suhara et al., 2022) and Sherlock (Hulsebos et al., 2019) on this task. Since Sherlock and DoDuo are designed for column annotation, we use the out-of-the-box model provided by the original teams. We restrict both to the fifteen target classes by setting the probabilities of non-target classes to zero. For DoDuo-Wiki, which supports a distinct set of classes, we perform a manual mapping to the class names used by DoDuo-VizNet and Sherlock. For TaBERT we train an additional classification layer on top of the pre-trained embeddings that these frameworks provide. We fix the embeddings to their pretrained values and learn the weights of the classification layer using five-fold cross-validation.
Results
Table 4 contains the results for the VizNet dataset. Our fm-based approach improves performance on the measured metrics of -score, precision and recall. The best performing method is Sherlock, narrowly beating DoDuo-VizNet, with a 0.954 score. If we consider methods which are not specifically pretrained on VizNet (note, which is also the test set) chorus is the best performing on all three metrics. It has comparable and precision to Sherlock, but percentage points lower recall.
Note in particular DoDuo-Wiki, which does not have access to VizNet at pretraining time, has a large regression in performance compared to DoDuo-Viznet, nearly half points. This drop is in line with previous results, see Section 5. We sanity-check the low scores of TaBERT by replicating previously reported scores (Koleva et al., 2021).
Chorus achieves a competitive throughput of 41 columns per second (col/s), comparable to Sherlock’s 50 col/s and exceeding DoDuo’s 7.3 col/s and TaBERT’s 4.5 col/s. This corresponds to benchmark completion in 13 minutes, as contrasted with over 2 hours 10 minutes for TaBERT. The average cost of GPT-3.5 calls for this task was 1.3¢ cents per 100 columns.
| -score | Precision | Recall | |
| DoDuo-VizNet | 0.876 | 89.4% | 87.2% |
| Sherlock | 0.954 | 96.2% | 94.6% |
| TaBERT | 0.321 | 32.6% | 32.0% |
| DoDuo-Wiki | 0.440 | 59.2% | 45.4% |
| Chorus | 0.891 | 91.2% | 88.8% |
4.3. Join-column prediction
Finally, we evaluate our approach’s ability to suggest which columns are the correct choice for a join, the join-column prediction task. We use the GitNotebooks dataset from (Yan and He, 2020), a collection of 4 million Python notebooks (and their associated relational tables) including 24 thousand joins collected from Github. One of the baselines, Trifacta Wrangler, requires manual execution and recording of each prediction. For that reason we restrict this benchmark to 300 randomly sampled tables.
Baselines
For this task, we compare with three baselines. Jaccard similarity, , is the first. Two columns are selected such that where . This is a commonly used approach in the literature (Chen et al., 2014; Dasu et al., 2002; Nargesian et al., 2023; Yan and He, 2020). Another baseline is Levenshtein distance (Levenshtein, 1966), which selects the pair of column names with the smallest edit distance between them. The final baseline is Trifacta Wrangler (Trifacta, 2023), a commercial product spun off from the Wrangler research line (Kandel et al., 2011). When joining two tables in this product, it suggests the keys on which to join them. As no api was available, we obtain all Trifacta predictions by joining manually.
Results
Table 5 shows the quality of estimates for our approach and the baselines. We measure the quality of the predictions by the same criteria as the previous tasks. By these metrics, our approach improves the quality of predictions and beats the next-best approach by a clear margin: score is improved by , precision by percentage points and recall by percentage points. This performance is maintained when scaling to the full dataset. On this task, our system has an average throughput of 23.5 predictions per second and cost approximately 5¢ cents per 100 predicted joins.
| -score | Precision | Recall | |
| Manually-run subset, | |||
| Jaccard | 0.575 | 60.7% | 54.7% |
| Levenshtein | 0.718 | 72.3% | 71.3% |
| Trifacta Wrangler | 0.823 | 82.6% | 82.0% |
| Chorus | 0.895 | 91.0% | 88.0% |
| Full dataset, | |||
| Jaccard | 0.458 | 63.3% | 35.9% |
| Levenshtein | 0.777 | 78.7% | 76.8% |
| Trifacta Wrangler | No api | ||
| Chorus | 0.912 | 93.2% | 89.4% |
4.4. Dataset contamination
Here we perform an experiment to validate whether any of the testing data occurred in the training corpus of the large-language model, an issue called dataset contamination or data leakage. Because these models are trained on internet data (Gao et al., 2021) and we use public benchmarks, they may have seen the test data in training.
We test on seven guaranteed-unseen tables (listed in the technical report (Kayali et al., 2023)) and their columns, all uploaded between April–June 2023 to the federal data repository Data.gov. They are guaranteed-unseen because the foundation model training was completed on or before March 2023. Repeating the supervised column-type annotation task as in Section 4.2, we measure a 0.857 score, 90.0% precision and 81.8% recall. This is within points, 0.1% precision and 5% recall of the benchmark results. See Table 6. The recall falling more is expected, as this reflects the datasets being more diverse and therefore difficult to classify.
| Dataset | -score | Precision | Recall |
| Public benchmark (VizNet) | 0.865 | 90.1% | 86.7% |
| Guaranteed-unseen | 0.857 | 90.0% | 81.8% |
4.5. System characteristics
Determinism
We examine the impact of nondeterminism in the foundation model on the performance of chorus. The randomness of the generation is controlled by the temperature hyperparameter. To assure that the results of chorus are reliable, we conduct the following experiment: we run the T2D table-class detection benchmark 25 times, five trials for each value of between . Figure 5 shows the result. Chorus’s performance is consistent: at the ideal temperature setting the score sees error bars of 0.01 points. The best performance is obtained at lowest temperature, —this is in contrast to NLP tasks like summarization that benefit from higher temperatures. In the prior experiments we use the default temperature, as to get chorus running with minimal hyperparameter tuning.
Alternative models
To demonstrate the versatility of this approach, we run chorus with three alternative, open-source foundation models on the table-class detection task. We consider Vicuna (Zheng et al., 2023), a variant of LLaMA (Touvron et al., 2023a) at two sizes: 13 billion parameters and 7 billion parameters. The more advanced model is LLaMA 2 (Touvron et al., 2023b), the SOTA open-source model with 70 billion parameters.
Table 7 shows the results. While OpenAI’s GPT model performs best, the best open-source model is very competitive. LLaMA 2 outperforms the best baseline model for this task—DoDuo-Wiki—by 0.136 points, on precision by 13.6 percentage points and on recall by 9.6 percentage points. This model lags behind the proprietary and larger GPT model by only a modest 0.03 points. Open-source LLMs are now compelling alternatives on the tested task.
| Table-class correctness | |||
| Model choice | -score | Precision | Recall |
| GPT-3.5 (175B) | 0.926 | 96.1% | 92.4% |
| LLaMA 2 (70B) | 0.893 | 92.2% | 86.5% |
| Vicuna/LLaMA (13B) | 0.713 | 79.2% | 64.1% |
| Vicuna/LLaMA (7B) | 0.713 | 75.3% | 67.5% |
Ablations
We conduct ablation experiments to measure the contribution of individual components of chorus. We remove one component at a time and note the loss of scores compared to the unaltered model. Figure 6 shows the results. First, we remove the demonstration from the prompt. This results in an F1 score loss of 0.03, a recall loss of 4.7 and a precision loss of 4.7 percentage points. Next, we remove the metadata where it is available. This results in a cumulative F1 score loss of 0.04, a recall loss of 5.1 and a precision loss of 5.6 percentage points. After that, we disable anchoring. This results in a cumulative F1 score loss of 0.389, a recall loss of 47.4 and a precision loss of 31.9 percentage points. From this the prevalence of hallucinations can be gleaned: the substantive score loss implies that hallucination is highly prevalent. The score improvement with anchoring implies that it can be recovered from: the fm mistakes are not severe but rather resemble the ground truth. Finally, we remove the prefixes from the prompt. This results in a cumulative F1 score loss of 0.736, a recall loss of 92.3 and a precision loss of 53.1 percentage points.
5. Discussion
Training data collection
A major advantage of a foundation-model approach is that there is no need for training on specific tasks. In contrast, TaBERT requires 26 million tables for training its embeddings. In (Deng et al., 2020), the use of 250 labels for one task is considered a “small dataset” by the authors and leads to subpar performance. In contrast, our prompts in Figure 3 use zero or none examples for each task.
Out-of-domain performance
We note a troubling pattern of a lack of cross-domain generalization in representation-learning approaches. The tested baselines degrade when used to embed tables not from the dataset the embeddings were trained on. This finding is in-line with prior work: regressions of up to 0.40 and 0.30 points when generalizing to new datasets have been reported (Deng et al., 2020; Hulsebos et al., 2021).
Flexibility
Another advantage of chorus we observe in the experiments is task adaptability. In the 1⃝ table-class detection task, we are able to switch the prediction domain easily. Restricting to the 33 classes used by the benchmark can be done by providing the permitted classes to the foundation model; allowing the model to generalize to other DBPedia classes (the unsupervised heading of Section 4.1) is as simple as omitting those instructions. Contextual information, such as table title or url, could be as easily added. Previously, such modifications would require retraining the embeddings.
Limitations and risks
We control the risk of dataset contamination by testing for it in Section 4.4. The performance of chorus on guaranteed-unseen datasets is comparable to those in public benchmarks, so good performance on the those benchmarks cannot be explained away as simple data contamination. Separately, formal linguistic fluency means that errors may fool human reviewers (Langer et al., 1978; Langford, 1999). This has been called subtle misinformation in prior work (Rae et al., 2021). Finally, prompts may not be robust to changes (Zhao et al., 2021).
Pattern-matching hypothesis
We posit that llms outperform representation learning approaches because they are able to extract a large number of linear transformations over the embedding space from their training data and compose them in simple ways. In the full technical report, we conduct an experiment contrasting llm activations with doduo weights, supporting this hypothesis (Kayali et al., 2023). Prior work (Olsson et al., 2022) supports this hypothesis and the limits of llms in composing these patterns, generally failing to combine patterns after about four reasoning steps (Dziri et al., 2023).
Future directions: Additional tasks. The above hypothesis suggests the promising performance may extend to many more tasks. Related tasks to be explored include schema auto-completion (Cafarella et al., 2008), where missing parts of a partial schema are suggested to the user; join-graph traversal, where successive tables to join on are suggested (Dong et al., 2022); and outlier detection, where erroneous data are detected. These are all promising because they involve composing simple patterns that are prevalent in fm’s training data. On the other hand, novel approaches will likely be needed to apply fms to tasks like data provenance (Green et al., 2007), since these involving tracking more patterns than fm’s are thought to currently support. Incorporating proprietary information may also be difficult to due the dearth of these patterns in the training data.
Private or domain-specific datasets
As with all the tested baselines, the foundation models are trained on public data. The distribution of data in the public sphere differs significantly from that in specialized domains or private data. It is worth investigating whether the observed capabilities continue to hold on e.g. enterprise data lakes. Further application to domain-specific ontology such as dron, a pharmaceutical ontology of drugs, would also be interesting to investigate.
Fine-tuned or distilled models
Larger models are of interest because if scaling laws continue to hold, their performance should improve (Wei et al., 2022b). The large inference cost of foundation models has sparked interest in model distillation: reducing the size of the model without incurring a reduction in specific capabilities (West et al., 2022; Dettmers et al., 2023) but at cost of degraded emergent abilities (Gudibande et al., 2023).
6. Related Work